MalayaJanae:

Which of the values in the set {2, 3, 4, 5} is a solution to the equation 2x + 4 = 10? (4 points) 2 3 4 5

AZ:

There's two ways we can solve your question: 1. We can plug in the answer choices into x and see which one makes the equation true or 2. We can solve the equation 2x + 4 = 10 for x For the first option, I'll give you an example We have 2, 3, 4, or 5. Let's test the number 2 and see if it is a solution to our equation. We have to replace the x with 2. 2(2) + 4 = 10 4 + 4 = 10 8 = 10 2 is not a solution because \(8 \neq 10\) A number that gives us 10 = 10 would be the solution Can you plug in the other numbers and see which one works?

MalayaJanae:

2, 3, 4, 5. Is the answer choice.

AZ:

If you would like to instead solve the equation, there are a few basic rules you have to keep in mind: When you're solving an equation, you're trying to isolate the variable (which is `x` in your question) Do you remember PEMDAS? To isolate the variable, you have to follow PEMDAS backwards. This means that you do the addition/subtraction first, and then you can do your multiplication/division. Now the key step is that whatever you're doing to your variable, you have to do the opposite of that on both sides. I'll give you an example that ties all of what I said together. 5x - 10 = 20 We first have to do the addition/subtraction We're subtracting 10 from 5x. To get our x all by itself we have to do the opposite on both sides. That means adding 10. 5x - 10 + 10 = 20 + 10 5x = 30 Now you can see we got rid of the 10 but we still haven't completely isolated 'x' 5x is x multiplied by 5. The opposite of multiplication is division so we have to divide 5 on both sides. 5x/5 = 30/5 x = 6

AZ:

I can see the answer choices as well but if I give you the answer, I would be doing you a great injustice. I can help you get to the answer (as I've done with my explanations and examples) but is there something that I can perhaps clarify further for you? What part of it do you not understand?

MalayaJanae:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ I can see the answer choices as well but if I give you the answer, I would be doing you a great injustice. I can help you get to the answer (as I've done with my explanations and examples) but is there something that I can perhaps clarify further for you? What part of it do you not understand? \(\color{#0cbb34}{\text{End of Quote}}\) Oh no! I understand it and love the help! Thank you!!!!!

AZ:

My pleasure!